### v1

# Expert Study on Interpretable Machine Learning Models with Missing Data

Abstract
Inherently interpretable machine learning (IML) models provide valuable insights for clinical decision-making but face challenges when features have missing values. Classical solutions like imputation or excluding incomplete records are often unsuitable in applications where values are missing at test time. In this work, we conducted a survey with 71 clinicians from 29 trauma centers across France, including 20 complete responses to study the interaction between medical professionals and IML applied to data with missing values. This provided valuable insights into how missing data is interpreted in clinical machine learning. We used the prediction of hemorrhagic shock as a concrete example to gauge the willingness and readiness of the participants to adopt IML models from three classes of methods. Our findings show that, while clinicians value interpretability and are familiar with common IML methods, classical imputation techniques often misalign with their intuition, and that models that natively handle missing values are preferred. These results emphasize the need to integrate clinical intuition into future IML models for better human-computer interaction.
keywords:
Interpretable machine learning, missing values, evaluation with clinicians
1 Introduction
Interpretable machine learning (IML) models have emerged as vital tools in high-stakes decision-making areas, such as healthcare (Ustun and Rudin, 2019a), sustainability (Jin et al., 2022), and criminal justice (Wang et al., 2023; Zhang et al., 2022), where transparency and accountability are paramount. Inherently interpretable models, such as linear or rule-based models, are well studied due to their apparent simplicity (Molnar, 2020; Rudin, 2019). While these models prioritize transparency, more advanced methods like interpretable decision sets (Rudin et al., 2022), checklists (Zhang et al., 2022) or falling rule lists (Wang and Rudin, 2015), effectively balance interpretability with predictive power, making them ideal for complex, real-world applications (Rudin et al., 2022). These models not only provide accurate predictions but also foster a deeper understanding of the model’s inner workings, enabling domain experts to make informed decisions (Molnar, 2020). Despite advancements in IML, aligning these models with clinical needs remains challenging (Rolnick et al., 2024). A key issue in healthcare applications is the prevalence of missing values both during training and in deployment, which can arise from incomplete patient records or errors during data collection (Salgado et al., 2016). Missing values are often addressed using imputation methods (Little and Rubin, 2019), which is general and powerful, but not always optimal under test-time missingness (Le Morvan et al., 2020b). Other methods like complete case analysis are inefficient (Penny and Atkinson, 2012), biased under non-uniform missingness, and are inapplicable to test-time cases with missing values. On the other hand, approaches that use simple imputations and missingness indicators (Le Morvan et al., 2020a) complicate interpretability. In short, missing values tend to make interpreting model predictions more difficult, potentially leading to issues during deployment.
To bridge the gap between IML model development and clinical use with test-time missingness, understanding clinicians’ needs is essential. A human-centered approach (Doshi-Velez and Kim, 2017) can guide model design to align with clinical expectations and improve healthcare reliability. We surveyed 71 clinicians to: 1) understand their interactions with predictive models, especially with missing values; 2) evaluate IML preferences aligned with clinical intuition; and 3) recommend improvements for handling missing data.
2 Methods
Participant recruitment
We conducted an online survey via LimeSurvey with 71 clinicians in July and August 2024 to evaluate their understanding of IML models and their effectiveness in handling missing data. Participants were recruited via email from TraumaBase network hospitals, and informed consent was obtained from all.
Experimental design
The survey, developed in English with input from trauma physicians and ML researchers, was translated into French by native speakers to ensure accuracy and cultural relevance. It included 41 questions on demographics, structured choice, and ranking tasks, Likert-scale statements, and open-ended questions, with both language versions provided in Appendix A. We conducted a pilot with five clinicians and incorporated their feedback. The survey aimed to understand clinicians’ interaction with IML when facing missing values at test time, using hemorrhagic shock prediction as an ICU decision-making example (Dutton et al., 2010) (details in Appendix A.2). The IML models included well-known interpretable classes such as risk scores (Ustun and Rudin, 2019b), linear model coefficients (Hastie et al., 2009), and decision trees (Breiman, 2017), all with comparable predictive performance (Molnar, 2020; Stiglic et al., 2020), as predictive performance was not the focus here. We also examined how different imputation methods influenced clinicians’ decisions. Presented imputation methods included native methods (e.g., XGBoost (Chen and Guestrin, 2016)) without imputation, implicit imputation allowing clinicians to adjust based on other feature values and judgment, explicit methods like mean or zero imputation using statistical replacements, and complex approaches such as MICE (Van Buuren and Groothuis-Oudshoorn, 2011) representing black-box methods. Decision trees handle missing data through surrogate splits. Figure LABEL:fig:IML_survey illustrates the survey interface, showing a case with one missing variable.
[…]
fig:IML_survey Refer to caption
Figure 1: Clinicians were given a patient sample (top) and three interpretable ML models (middle: decision tree, logistic regression; bottom: risk score) predicting hemorrhagic shock. We assessed their interaction with the models with missing values.
Data analysis
We analyzed multiple-choice responses using descriptive statistics, while Likert scale and categorical data were examined with factor analysis for mixed data (FAMD) using the FactoMineR package in R (Lê et al., 2008) to identify patterns and group perspectives. Free-text responses were analyzed for central tendency and variability, with selected French responses translated into English. Figure LABEL:fig:survey_example in the Appendix shows the experimental setup.
3 Results
3.1 Sample characteristics
The survey was administered to 214 clinicians across 29 trauma centers in France, part of the TraumaBase network, which collects data on over 50,000 trauma cases nationwide. We received 55 responses, with 20 (36.4%) fully completed and 35 (63.6%) partially completed. All responses were analyzed, so answer counts vary by question. The average age was 41.2 years, with a gender distribution of 22 females and 33 males; most participants were clinical doctors (83%) in Anesthesiology/Intensive Care (84%) with medical degrees (84%). Participant demographics are in Table LABEL:app:tab:stats_participants. Our FAMD analysis identified four clusters of clinicians based on AI/ML attitudes, familiarity with IML, and missing value handling preferences (analysis in Appendix A.4). The most prominent cluster (25%) comprises frequent AI users, highly familiar with IML models, and prefer implicit imputation over black-box methods.
Current practices with missing values
Clinicians use various strategies to manage missing data in clinical scores, primarily relying on clinical intuition or other available clinical information. Many mentioned using plausible values based on clinical context or the patient’s history to estimate missing values. Some prefer worst-case imputation to avoid underestimating risk, though one noted it may lead to overtraining and resource inefficiency. Others choose to partially calculate or omit scores if key data is missing, feeling that imputation could ”skew the score concept.” One clinician also suggested using alternatives, like thromboelastography when prothrombin time is unavailable.
3.2 Evaluation of IML with missing values
Our survey focuses on evaluating how clinicians interact with tools, investigating how different predictive models handle (or natively handle) missing values, and how clinicians interpret these models to make informed decisions. We also examine what clinicians find lacking in current IML approaches and what improvements are needed to make these models more useful for clinical decision-making.
Understanding approaches to handling missingness in prediction tasks
To understand clinicians’ perspectives on common imputation methods, we first provided average hemoglobin values and tested reactions to zero-imputed data (Table LABEL:tab:Imputation_ranking). Most clinicians reported that average values were irrelevant for decision-making, emphasizing the need for individualized consideration, as one explained: “Each patient is unique, and I don’t see the relevance of giving an average value.” Additionally, some clinicians expressed concern that mean imputation might underestimate patient risk, potentially minimizing preparedness for emergency scenarios. When presented with a zero-imputed value, many clinicians indicated they would ignore the feature, though some noted the importance of clearly marking any imputation along with associated uncertainty. For decision trees, using surrogate values was explored, but this method was less familiar to clinicians and caused some confusion. However, many clinicians noted that its usefulness depends on clearly indicating where the substitution occurs. More complex methods as MICE which frequently used and cited in medical journals (e.g.,  Austin et al. (2021); Hegde et al. (2019); Luo (2022)) as an imputation tool for clinical data. While clinicians recognize MICE for its effectiveness in imputing missing values, it does not necessarily enhance predictive performance. MICE is designed for inference, not primarily for supervised learning. In our survey, black-box methods like MICE were selected in less than 10% of cases (6/51), indicating that clinicians hesitate to use methods they cannot interpret easily. MICE’s low preference may also stem from it not being commonly taught to physicians.
Clinicians are confident but not always correct under missing values
Clinicians showed high confidence (9-10 out of 20 answers) in explaining risk scores and decision trees, even with missing values, and interpreted the model outcome correctly. For logistic regression, however, confidence was lower and accompanied by uncertainty; only 6 participants were correct, with 4 out of 6 correct among those expressing confidence in its usefulness with missing values.
Requirements to support interaction between (future) IML & clinicans under missing values
Four out of 12 clinicians emphasized the need for clear guidelines and transparency in handling missing data, with one suggesting the presentation of multiple imputed values to allow intuitive selection along with communicating uncertainty. The reliability of models for managing missing data were seen as essential, with two clinicians stressing that ML conclusions should be accessible to non-technical users, recommending external validation and visual aids to clarify the impact of imputation. Clinicians underscored the importance of integrating clinical intuition with ML outputs, emphasizing that “the human element and intuition based on experience and the CLINICAL EXAMINATION OF THE PATIENT!” should complement, not replace, the model. Their preference for implicit over explicit imputation methods reflects a focus on transparency and alignment with clinical reasoning, as implicit methods enhance situational awareness.
4 Discussion
Context
While Janssen et al. (2010) argues that simple methods for handling missing data can lead to misleading results, others advocate for multiple imputation techniques as they enhance validity in clinical research (Austin et al., 2021). Following Tonekaboni et al. (2019), who emphasized evaluating clinical explainability in high-stakes settings, IML models must undergo rigorous validation to assess performance, interpretability, and legal implications (Wiens et al., 2019). Clinicians can build trust in IML models by understanding them and through positive experiences where model outputs align with domain knowledge, enhancing appreciation for the model’s utility while ensuring accountability for patient outcomes (Kelly et al., 2019). Randomized controlled trials (RCTs) can validate models, but are costly, time-consuming, and may lack generalizability across diverse clinical settings (Shortliffe and Sepúlveda, 2018). Successful deployment involves robust model development, presenting uncertainties (Doshi-Velez and Kim, 2017), and training clinicians on the impact of missing values to integrate these tools into workflows (Topol, 2019; Maddox et al., 2019).
Limitations
Our survey focused on ICU and trauma specialists with foundational knowledge of clinical ML systems. However, it may reflect a self-selection bias due to the nature of participant recruitment. Focused on IML models for prediction with missing values, the findings may not directly apply to other paradigms, like reinforcement learning, or to settings like outpatient care. Only models from three IML classes were explored and the scope was limited to clinicians in France, highlighting the need for broader research across different countries and healthcare systems to generalize and externally validate the results.
Development of IML models with missing values: Interpretation of clinical needs
Our results reveal that current models often fall short of supporting clinical decision-making when faced with missing data at test time. To improve adoption by healthcare professionals, future IML models should better align with clinical intuition and incorporate clinical experience when handling missing data.
While imputation methods like zero, mean, or black-box imputation may facilitate model training, they are less favored for test-time use. Worst-case imputation, although cautiously applied by some clinicians, can introduce statistical bias, skew predictions, and misrepresent patient risk. Clinicians prefer that imputation, if necessary, incorporate patient-specific data and clinical experience, rather than generic values like zeros or cohort means. Moreover, any imputation used should be transparent, as clinicians bear responsibility for decisions impacting patient outcomes. Future methods should either avoid reliance on imputation or employ more transparent imputation strategies that are suitable both during training and in use. Models that natively support missing values, like those by Chen and Guestrin (2016) and Stempfle and Johansson (2024), eliminate or reduce the need for imputation during training or testing. Alternatively, more transparent imputation methods can enhance trust. For example, recent work by Chen et al. (2023) on explainable boosting machines provides insights into missing data mechanisms, helping to understand causes and mitigate risks. As mentioned by several respondents, future models should also respect clinical reasoning by integrating patient history and the sequential nature of medical information. Current IML methods often overlook this aspect, making alternatives like reinforcement learning potentially more suited to capturing the complexity of clinical decision-making.
Conclusion
The results reveal that clinicians value IML models but find traditional approaches like imputation inadequate when data is missing at test time. They prefer models that can natively handle missing values, as imputation often conflicts with clinical intuition. This underscores the need for ML models aligned with clinical needs, particularly in high-stakes areas like hemorrhagic shock prediction. The study highlights clinicians’ preferences for advanced methods to enhance decision-making and patient outcomes with more reliable, interpretable tools.
[…]
We would like to express our great gratitude to the healthcare professionals who participated in the experiment. We appreciate that they voluntarily devoted their time to complete the experiment. This paper and the research behind it would not have been possible without the significant feedback they delivered.
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A
A.1 Survey setup
LimeSurvey is an open-source online survey tool that allows users to create, manage, and analyze surveys. It provides an easy-to-access interface where participants can complete surveys via a web link. For your usage, participating clinics could access LimeSurvey to fill out surveys or collect data in a structured format. The tool offers customizable survey templates, various question types, and features like branching logic to tailor the survey experience, making it user-friendly and efficient for collecting data from clinical settings.
Survey instrument
The survey was developed in English with input from trauma physicians and ML researchers and then translated into French by native speakers for accuracy and cultural relevance. The 41 questions included demographics, single/multiple-choice items with comments, ranking tasks, open-ended questions, and statement blocks on a 5-point Likert scale (Taherdoost, 2019). Likert scales gauged attitudes, choice questions collected structured data, ranking tasks prioritized preferences, and open-ended questions offered deeper insights. Appendix A.3 contains the questionnaire in both languages.
[…]
fig:survey_example Refer to caption
Figure 2: Experimental setup. Physicians are shown a data entry of a patient with 9 features in an interpretable machine-learning model along with the questions. After the data is gathered qualitative and quantitative methods are used to analyze the results.
A.2 Use case
We aim to understand the challenges of diagnosing hemorrhagic shock in trauma patients, a leading cause of early, preventable death in cases of severe injury (Kauvar and Wade, 2005; Dutton et al., 2010). Timely intervention is crucial, as delays can significantly worsen prognosis. For our survey, hemorrhagic shock is defined as a composite outcome that includes any patient who meets one or more of the following criteria: 1) administration of at least one packed red blood cell (RBC) in the resuscitation room, 2) transfusion of four or more RBC units within six hours of admission, 3) the need for a hemorrhage control procedure (such as interventional radiology or surgery) within six hours of admission, or 4) death from hemorrhagic shock within 24 hours (Hamada et al., 2018). The data in the use case include both continuous and categorical features, covering clinical variables like heart rate and non-clinical factors such as prior clinician decisions, including ordering specific lab tests.
A.3 Survey questions
The survey contained along with the demographic questions four domains. The domains are: Physicians’ Current Decision-Making Behavior in Diagnosing Hemorrhagic Shock in Trauma Patients (Current Decision-Making), Interpretation of ML models in the context of predicting with missing values at test time for a clinical use case (Interpretation of ML models and missing values), Preferences for using IML prediction model in daily workflows (Preferences in clinical care), Willingness to incorporate IML for prediction tasks in the context of missing values for diagnosing hemorrhagic shock and organ failure in clinical practice (Willingness)
In French: Les domaines sont les suivants: Comportement décisionnel actuel des médecins concernant le diagnostic du choc hémorragique chez les patients traumatisés (Opinion de l’Intelligence artificielle/Machine learning), Interprétation des modèles d’apprentissage automatique dans le contexte de la prédiction avec des valeurs manquantes au moment du test (Interprétation des modèles d’apprentissage automatique et des valeurs manquantes), Préférences pour l’utilisation du modèle de prédiction de l’IML dans les flux de travail quotidiens (Préférences de soins cliniques), Volonté d’incorporer l’IML pour les tâches de prédiction dans le contexte des valeurs manquantes pour le diagnostic du choc hémorragique et de la défaillance d’organe dans la pratique clinique (Volonté).
[…]
tab:rationales_english
Table 1: All questions in survey in the English language
Domain Question
Current Decision-Making Behavior Q: Are you currently relying on decision support tools, such as risk scores, for diagnosing hemorrhagic shock in trauma patients? If yes, which ones and how do they support you?
Q: Are you currently making use of any risk scores, e.g., RedFlag or ABC Score to support you in the diagnosis of hemorrhagic shock for trauma patients? If yes, please name the risk score and how you interpret the risk score when applying it to patients with missing values.
Q: For which group of patients is the diagnosis of hemorrhagic shock the most difficult?
Q: Which factors/features are most important when diagnosing hemorrhagic shock in trauma patients?
IML models and missing values at test time Q: Please respond to all of the questions in your own words and provide as much detail as possible to understand your reasoning. What would your prediction (or the range of the prediction) be?
Q: The average hemoglobin value in this data set for females is 12.70 g/dl. Does that knowledge affect your prediction assuming the patient is a woman?
Q: Assuming the prediction made for the patient (no. 32450) was made by adding a zero for the missing value, which resulted in a positive prediction indicating they will have hemorrhagic shock, is the representation of risk scores despite the missing value helpful for your decision-making?
Q: Do you trust the risk score? Motivate your answer briefly.
Q: Please respond to all of the questions in your own words and provide as much detail as possible to understand your reasoning. What would the prediction (or the range of the prediction) be?
Q: The average hemoglobin value in this data set for females is 12.70 g/dl. Does that knowledge affect your prediction assuming the patient is a woman?
Q: Assuming that the coefficients were derived from a dataset completed using mean imputation, a positive prediction was then made for the patient. Would you trust that the patient has a hemorrhagic shock?
Q: Do you trust the linear model? Motivate your answer briefly.
Q: Please respond to all of the questions in your own words and provide as much detail as possible to understand your reasoning. What would the prediction (or the range of the prediction) be?
Q: The average hemoglobin value in this data set for females is 12.70 g/dl. Does that knowledge affect your prediction assuming the patient is a woman?
Q: Assuming the prediction made for the patient (no. 32450) was made by replacing the missing value using a surrogate split, a common approach in decision trees to handle missing values, which resulted in a positive prediction indicating they will have hemorrhagic shock. In a surrogate split, an alternative value is used to decide whether to go left or right in the tree. Is the representation of risk scores despite the missing value helpful for your decision-making?
Q: Do you trust the decision tree? Motivate your answer briefly.
Preferences in clinical care Q: Missing values in decision trees and risk scores can be handled by implicit imputation, which means that a value for the missing one can be imputed by considering the values of the other features. This might result in individual values depending on the clinician. Rank which method—decision tree or risk score—you think implicit imputation is more suitable for?
Q: Another alternative is to use so-called black box models that perform the imputation by using methods that use complex, often non-transparent algorithms to fill in missing values in a dataset. Examples might be using Multivariate Imputation by Chained Equations (MICE). For which interpretable model would you most likely choose the black box imputation? Rank them according to your preference.
Q: Do you prefer implicit or explicit imputation and why? You can consider aspects such as ease of understanding, usefulness in communicating your prediction, or intuitiveness.
Q: Which missingness handling methods do you like best for each IML? Mark 3 in total with an x and add 0 in the other cells.
Willingness Q: What is lacking in interpretable ML models for effective clinical use when dealing with missing values?
[…]
tab:rationales_french
Table 2: Toutes les questions de l’enquête sont en français.
Domaine Question
Opinion de l’IA/ML Q: Utilisez vous actuellements des outils d’aide à la décision, tels que les scores de risque, pour diagnostiquer le choc hémorragique chez les patients traumatisés? Si oui, lesquels et comment vous aident-ils?
Q: Utilisez-vous actuellement des scores de risque, par exemple RedFlag ou ABC Score, pour vous aider à diagnostiquer le choc hémorragique chez les patients traumatisés ? Dans l’affirmative, veuillez indiquer le score de risque et comment vous l’interprétez lorsque vous l’appliquezà des patients dont les valeurs sont manquantes.
Q: Pour quel groupe de patients le diagnostic de choc hémorragique est-il le plus difficile?
Q: Quels facteurs/caractéristiques sont les plus importants lors du diagnostic du choc hémorragique chez les patients traumatisés?
Interprétation des modèles IML Q: Veuillez répondre à toutes les questions avec vos propres mots et fournir autant de détails que possible pour comprendre votre raisonnement. Quelle serait votre prédiction (ou la plage de prédiction)?
Q: La valeur moyenne d’hémoglobine dans cet ensemble de données pour les femmes est de 12,70 g/dl. Cette connaissance affecte-t-elle votre prédiction en supposant que le patient est une femme?
Q: En supposant que la prédiction faite pour le patient (n° 32450) a été réalisée en remplaçant la valeur manquante par un zéro, ce qui a abouti à une prédiction positive indiquant qu’il y aura un choc hémorragique, la représentation des scores de risque malgré la valeur manquante est-elle utile pour votre décision? Motivez brièvement votre réponse.
Q: Faites-vous confiance au score de risque? Motivez brièvement votre réponse.
Q: Veuillez répondre à toutes les questions avec vos propres mots et fournir autant de détails que possible pour comprendre votre raisonnement. Quelle serait la prédiction (ou la plage de prédiction)?
Q: La valeur moyenne d’hémoglobine dans cet ensemble de données pour les femmes est de 12,70 g/dl. Cette connaissance affecte-t-elle votre prédiction en supposant que le patient est une femme?
Q: En supposant que les coefficients étaient dérivés d’un jeu de données complété par imputation moyenne, une prédiction positive a alors été faite pour le patient. Seriez-vous convaincu que le patient présente un choc hémorragique?
Q: Faites-vous confiance au modèle linéaire?
Q: Veuillez répondre à toutes les questions avec vos propres mots et fournir autant de détails que possible pour comprendre votre raisonnement. Quelle serait la prédiction (ou la plage de prédiction)?
Q: La valeur moyenne d’hémoglobine dans cet ensemble de données pour les femmes est de 12,70 g/dl. Cette connaissance affecte-t-elle votre prédiction en supposant que le patient est une femme?
Q: En supposant que la prédiction pour le patient (no. 32450) a été faite en utilisant une technique appelée division de substitution pour traiter la valeur manquante, cette méthode est en effet couramment employée dans les arbres de décision. Dans une division de substitution, une autre valeur est utilisée pour décider du chemin dans l’arbre (gauche ou droite). Imaginons qu’une division de substitution ait été effectuée et qu’elle ait conduit à une prédiction positive indiquant un choc hémorragique. La représentation des arbres décisionnels avec la division de substitution pour traiter la valeur manquante vous aide-t-elle dans votre prise de décision?
Q: Faites-vous confiance à l’arbre de décision? Motivez brièvement votre réponse.
Préférences de soins cliniques Q: Les valeurs manquantes dans les arbres de décision et les scores de risque peuvent être traitées par imputation implicite, ce qui signifie qu’une valeur pour la valeur manquante peut être imputée en considérant les valeurs des autres caractéristiques. Cela peut entraîner des valeurs individuelles en fonction du clinicien. Classez quelle méthode (arbre de décision ou score de risque) pour laquelle vous pensez que l’imputation implicite est la plus adaptée?
Q: Une autre alternative consiste à utiliser des modèles dits de boîte noire qui effectuent l’imputation en utilisant algorithmes plus complexes, souvent non transparents, pour combler les valeurs manquantes dans un ensemble de données. Des exemples pourraient être l’utilisation de l’imputation multivariée par équations chaînées (MICE). Pour quel modèle interprétable choisiriez-vous le plus probablement l’imputation par boîte noire? Classez-les selon vos préférences.
Q: Préférez-vous l’imputation implicite ou explicite et pourquoi? Pouvez-vous prendre en compte des aspects tels que la facilité de compréhension, l’utilité de communiquer votre prédiction à un collègue ou l’intuitivité?
Q: Quelles méthodes de gestion des absences préférez-vous pour chaque modèle interprétable? Marquez vos 3 favoris d’un x et ajoutez 0 dans les cellules.
Volonté Q: Que manque-t-il aux modèles ML interprétables pour une utilisation clinique efficace lorsqu’il s’agit de valeurs manquantes?
A.4 Additional results
This section presents further findings, including figures on the attitudes of different cohorts and the characteristics of identified clusters within these cohorts. Additionally, we provide detailed rankings of participants’ preferences for various imputation methods across different IML techniques.
Cohort analysis
We conducted a Factor Analysis for Mixed Data (FAMD) Josse and Husson (2012), revealing four clusters among the survey responses. Missing values e.g. unanswered questions, were imputed using Principal Component Analysis (PCA) (Josse and Husson, 2012) for continuous variables and the most frequent category for categorical data. The primary goal FAMD Josse and Husson (2012) and the clustering is descriptive—to provide a multivariate view of clinician groups rather than relying on just one or two variables. We used FAMD Pagès (2014); Escofier and Pages (1994), a low-rank approximation method specifically suited for mixed data (both continuous and categorical), serving as the counterpart to PCA Greenacre et al. (2022) for such datasets. FAMD is particularly useful for analyzing questionnaire results, as it balances the influence of both quantitative and categorical variables while computing distances between observations (in this case, clinicians). The first principal component of FAMD is a quantitative variable most related to all others, measured by squared correlation for continuous variables and $R^{2}$ for categorical variables (similar to how PCA works for quantitative data , with $R^{2}$ analogous to an ANOVA model for categorical data).
We identified four clusters (Figure in Appendix LABEL:fig:cohort_cluster) based on clinicians’ attitudes towards AI/ML, their familiarity and trust in IML, and their preferences for handling missing values, regardless of demographic factors. Cluster 1 includes clinicians (16% of 55, 9 people) with low understanding and use of risk scores, little familiarity with decision trees, and dissatisfaction with all three models when dealing with missing values. Despite this, they are comfortable using black-box models like MICE but prefer implicit imputation and surrogate splits. Cluster 2 consists of clinicians (22%, 12 people) who show high levels of fear of AI, blind trust in its capabilities, and concerns about job loss, alongside perceiving current AI solutions as low-quality. They are familiar with decision trees and favor zero imputation or surrogate splits. Cluster 3 has substantial missing demographic data and may need to be excluded from the analysis (36%, 20 people). Cluster 4 comprises clinicians (25%, 14 people) who use AI daily, are highly familiar with all three IML models, and feel confident explaining these models with or without missing values. They favor implicit imputation over black-box models and surrogate splits.
[…]
app:tab:stats_participants Characteristic Total (N=55) Complete answers (N=20) Age in years, mean (SE) 41,18 (1,13) 39,35 (1,55) Female, n (%) 22 (40%) 7 (5%) Male, n (%) 33 (60%) 13 (65%) Highest received educational degree, n (%)    University bachelor/master/postgraduate education 9 (16,36%) 3 (15%)    Medical doctoral degree 37 (67,27) 12 (60%)    Associate Professor/Professor 9 (16,36%) 5 (25%) Current level of occupation, n (%)    Clinical doctor 46 (83,64%) 19 (95%)    Intern/Trainee/Non-fully graduated clinician 3 (5,45%) 0    Other 6 (10,91%) 1 (5%) Medical specialty, n (%)    Anesthesiology /// Intensive Care 46 (83,64%) 16 (80%)    Emergency medicine 5 (9,09%) 2 (10%)    Other 4 (7,27%) 2 (10%)
Table 3: Descriptive statistics of survey participants. Categorical variables have been simplified for brevity: The ’Other’ category for current occupation included roles such as nurse, clinical studies technician, hospital clinical trials coordinator, clinical manager assistant, and retired professional. For medical specialization, it included fields like rehabilitation, pediatrics, surgery, anesthesiology, bioinformatics, pediatric orthopedic surgery, and clinical research associate.
[…]
tab:Imputation_ranking
Table 4: The table presents the results of clinicians’ rankings of implicit and explicit imputation methods across different IML techniques. Clinicians were asked to evaluate and rank three IML methods, focusing on their preferences between implicit imputation, which considers other feature values for handling missing data and allows for expert adjustments, and explicit imputation, which uses methods like zero or mean imputation. The table also includes rankings for more complex and less transparent algorithms, such as Multivariate Imputation by Chained Equations (MICE) (Van Buuren and Groothuis-Oudshoorn, 2011), used to fill in missing data.
Rank which method—decision tree or risk score—you think implicit imputation is more suitable for?
Rank Risk Scores Decision Tree Sum
Rank 1 9 8 17
Rank 2 8 9 17
For which interpretable model would you most likely choose the black box imputation?
Rank Risk Scores Decision Tree Coefficients in Linear Models
Rank 1 4 4 7
Rank 2 3 5 7
Rank 3 6 4 5
Sum 13 13 19
[…]
fig:cohort_cluster Refer to caption
Figure 3: The cohort is divided into four clusters, each reflecting different attitudes towards AI/ML, varying levels of familiarity with IML, both within and outside the context of missing values and imputation preferences along with general trust in IML. In the color coding, blue indicates that the mean for this cluster is significantly lower than the global mean, while red indicates that the mean is significantly higher. White or light shades of blue and red suggest no significant difference within the group regarding the variable. Demographical features were only added to describe the clusters.